A cylinder of radius 4.09 cm and a spherical shell of radius 7.47 cm are rolling without slipping along the same floor. The two objects have the same mass. If they are to have the same total kinetic energy, what should the ratio of the cylinders angular speed to the spherical shells angular speed be? Number cyl sal
Show all work and provide the calculation
Answer #1

here,

radius of cyclinder,rc = 4.09 cm = 0.0409 m

radius of spherical shell , rs = 7.47 cm = 0.0747 m

the total kinetic energy of cyclinder , KEc = 0.5 * m * vc^2 + 0.5 * Ic* wc^2

KEc = 0.5 * m * (rc * wc)^2 + 0.5 * ( 0.5 * m * rc^2) * wc^2

KEc = 0.75 * m * ( rc*wc)^2

the total kinetic energy of sphereical shell, KEs = 0.5 * m * vs^2 + 0.5 * Is* ws^2

KEs = 0.5 * m * (rs * ws)^2 + 0.5 * ( 0.67 * m * rs^2) * ws^2

KEs = 0.83 * m * ( rs*ws)^2

as they have the same total kinetic energy

KEc = KE

0.75 * m * ( rc*wc)^2 = 0.83 * m * ( rs*ws)^2

0.75 * ( 0.0409 *wc)^2 = 0.83 * ( 0.0747*ws)^2

wc /ws = 1.92